# BECS: A Privacy-Preserving Computing Sharing Mechanism for 6G Computing Power Network

### v2

@@ VII-B Simulation of Proposed Computing Allocation Scheme
In this part, we compare NSGA-III, MOEA/D, NSGA-II-SDR, and NSGA-III-KDR in optimizing the proposed computing allocation MOOP. We initialize the population with 50% of the computing devices randomly occupied, then perform the optimization using each of the four algorithms separately. We calculate the changes in computing resources between the final and initial populations, while ensuring algorithmic convergence. Due to the uncertainty of the evolutionary process, besides comparing the overall data (represented by a box plot containing five horizontal lines representing the maximum, upper quartile, median, lower quartile, and minimum value from top to bottom) and the optimal data based on TOPSIS (represented by triangles), the average data (represented by squares) are also compared to comprehensively assess the performance of the four algorithms.
@@
As resource utilization improved, Fig. 8 to Fig. 8 illustrate the changes in the other five objectives considered by the computing allocation MOOP. Except for load balancing, improvements in all other objectives are observed with the optimization based on NSGA-III-KDR, attributed to enhanced resource utilization. The superior diversity of NSGA-III-KDR facilitates a more even distribution of computing resources, as particularly evidenced by the significant increase in privacy entropy. Although NSGA-III exhibits similar trends, it is less effective than NSGA-III-KDR. Notably, with NSGA-III-KDR and NSGA-III, increased resource utilization results in higher delays and sharing revenue, while energy consumption decreases as more tasks are offloaded to $\mathbb{U}$, leveraging all available resources within 6G CPN comprehensively. The optimization strategy of NSGA-II-SDR favors offloading tasks to $\mathbb{E}$ and $\mathbb{C}$, significantly reduces delay, but uniquely results in positive growth in load balancing among the four algorithms. The optimal solution of MOEA/D surpasses all other algorithms, achieving the highest resource utilization and substantial improvements in all objectives, except for load balancing, and the decline in load balancing suggests a rational allocation of computing resources. However, the overall data of MOEA/D is the worst, with a negative optimization of 12.99%. Furthermore, although its median or average increase in resource utilization is less than that of NSGA-III-KDR, the more substantial improvement in load balancing suggests a less favorable overall evolution. Since the selection of the optimal solution using TOPSIS is random, not every iteration of MOEA/D yields a solution that outperforms those of other algorithms.
@@ VII-C Simulation of Proposed Pseudonymous Scheme
TABLE IV: Execution time of cryptographic operations
@@
In this part, we evaluate the computational overhead of the proposed pseudonym scheme by comparing it with several existing pseudonymous schemes. Each operation and phase are tested 20 times separately, and the average value is recorded as the experimental data to eliminate the influence of hardware fluctuations during the operation. First, the BN curve [44] with a 128-bit security level is selected to implement the bilinear group. The execution time of basic cryptographic operations are shown in Table IV. Generally, the execution times required for bitwise XOR and modular multiplication are significantly lower compared to other cryptographic operations, and can thus be disregarded. Moreover, we utilize SHA-256 as the hash function. Then, we assess the time costs associated with system initialization and pseudonym generation (SIPG), certificate issuance or message signing(CIMS), and identity or message verification (IDMV) in the pseudonym authentication process, comparing the proposed scheme with those in [45, 46, 47, 48], as shown in Table V.
TABLE V: Comparison of computational overhead
Scheme SIPG CIMS IDMV Total
Bagga et al.’s scheme [45] $12.8674T_{pm}^{ecc}+2T_{mtp}+1T_{ha}=$ 12.867 $ms$ $1.863T_{pm}^{ecc}+3T_{pa}^{ecc}+3T_{ha}=$ 1.86 $ms$ $6.8473T_{bp}+4T_{pm}^{ecc}+4T_{pa}^{ecc}+3T_{ha}=$ 6.847 $ms$ $21.574$ $ms$
Shen et al.’s scheme [46] $9.5053T_{bp}+4T_{pm}^{bp}+1T_{pa}^{bp}+2T_{ep}+4T_{ha}=$ 9.505 $ms$ $2.7451T_{bp}+1T_{pm}^{bp}+1T_{ep}+1T_{ha}=$ 2.745 $ms$ $5.4812T_{bp}+2T_{pm}^{bp}+1T_{pa}^{bp}+1T_{ep}+1T_{ha}=$ 5.481 $ms$ $17.731$ $ms$
Yang et al.’s scheme [47] $10.52217T_{pm}^{ecc}+5T_{pa}^{ecc}+11T_{ha}=$ 10.522 $ms$ $0.0011T_{ha}=$ 0.001 $ms$ $1.863T_{pm}^{ecc}+4T_{pa}^{ecc}+2T_{ha}=$ 1.86 $ms$ $12.383$ $ms$
Wang et al.’s scheme [48] $2.4754T_{pm}^{ecc}+1T_{pa}^{ecc}+2T_{ha}=$ 2.475 $ms$ $5.5719T_{pm}^{ecc}+4T_{pa}^{ecc}+5T_{ha}=$ 5.571 $ms$ $1.863T_{pm}^{ecc}+2T_{pa}^{ecc}+4T_{ha}=$ 1.86 $ms$ $9.906$ $ms$
Our proposed scheme $4.3317T_{pm}^{ecc}+1T_{pa}^{ecc}+4T_{ha}=$ 4.331 $ms$ $1.8573T_{pm}^{ecc}+1T_{pa}^{ecc}+2T_{ha}=$ 1.857 $ms$ $2.4764T_{pm}^{ecc}+2T_{pa}^{ecc}+2T_{ha}=$ 2.476 $ms$ $8.664$ $ms$
Next, considering the diversity of computational devices, we test the portability of the proposed pseudonym scheme. We consider five different devices, including a workstation with an Intel i9-12900k@3.9GHz and 64GB RAM (Intel i9), a computer with an i7-12700@2.1GHz and 32GB RAM (Intel i7), a computer with an i5-8500@3GHz and 8GB RAM (Intel i5), a smartphone with Snapdragon 7+ Gen 2 and 16GB RAM (Snapdragon 7+ Gen 2), and a MacBook with an M1 chip and 16GB RAM (Apple M1). We test the time consumption for each of the five phases of the pseudonym authentication process on these devices: System Initialization (SI), Computing Devices Registration (CR), Pseudonym Generation (PG), Certificate Issuance (CI), and Identity Verification (IV), including the Total Time (TT) for the entire process, as shown in Fig. 9. Time consumption varies across devices due to differences in CPU performance. Specifically, the MacBook requires the longest total time to complete a pseudonym authentication, only $20.813\ ms$. The SI is the most time-consuming, as it involves generating system keys and registering user information. However, typically, SI is only performed once, whereas CR, PG, CI, and IV take less than $8\ ms$ across all devices, with PG taking the longest at $7.628\ ms$ on the smartphone. This demonstrates the good portability and lightweight of the proposed pseudonym scheme.
@@ VIII Conclusion
This article investigates improving computing resource utilization within 6G CPN and proposes BECS, a privacy-preserving computing sharing mechanism. We consider various communication, computing, and service factors in 6G CPN, modeling them as a five-objective MOOP. Meanwhile, we utilize the proposed NSGA-III-KDR to find optimal solutions for this MOOP. Additionally, we introduce a novel pseudonym scheme to protect the privacy of users engaged in computing sharing. Extensive simulations demonstrate the effectiveness of BECS. Moving forward, we intend to further explore computing sharing in dynamic real-time scenarios and address security and trustworthiness issues in task offloading.

### v3

@@ VII-B Simulation of Proposed Computing Allocation Scheme
In this part, we compare NSGA-III, MOEA/D, NSGA-II-SDR, and NSGA-III-KDR in optimizing the proposed computing allocation MOOP. Specifically, each chromosome in the evolutionary algorithm represents a candidate computing resource allocation scheme for the entire network, where each gene corresponds to the binary occupancy status of a specific computing device. Consequently, the population constitutes a diverse set of potential allocation schemes. To simulate a representative snapshot of the dynamic CPN environment, we initialize the population with 50% of the computing devices randomly occupied, then perform the optimization using each of the four algorithms separately. This simulation evaluates the algorithm’s ability to obtain a high-quality allocation solution within a single decision epoch, demonstrating its effectiveness in adapting to the heterogeneous and resource-constrained conditions of that specific moment. We calculate the changes in computing resources between the final and initial populations, while ensuring algorithmic convergence. Due to the uncertainty of the evolutionary process, we evaluate the performance of the four algorithms using three statistical indicators. Specifically, the overall distribution is presented via box plots (showing the maximum, upper quartile, median, lower quartile, and minimum), the optimal values based on the TOPSIS method are denoted by triangles, and the average values are represented by squares. This multi-dimensional comparison enables a more comprehensive assessment of algorithm performance.
@@
As resource utilization improved, Fig. 8 to Fig. 8 illustrate the changes in the other five objectives considered by the computing allocation MOOP. Except for load balancing, improvements in all other objectives are observed with the optimization based on NSGA-III-KDR, attributed to enhanced resource utilization. The superior diversity of NSGA-III-KDR facilitates a more even distribution of computing resources, as particularly evidenced by the significant increase in privacy entropy. Although NSGA-III exhibits similar trends, it is less effective than NSGA-III-KDR. Notably, with NSGA-III-KDR and NSGA-III, increased resource utilization results in higher delays and sharing revenue. Meanwhile, energy consumption decreases as more tasks are offloaded to $\mathbb{U}$, leveraging all available resources within 6G CPN comprehensively. The optimization strategy of NSGA-II-SDR favors offloading tasks to $\mathbb{E}$ and $\mathbb{C}$, significantly reducing delay, but uniquely resulting in positive growth in load balancing among the four algorithms. The optimal solution of MOEA/D surpasses all other algorithms, achieving the highest resource utilization and substantial improvements in all objectives, except for load balancing, and the decline in load balancing suggests a rational allocation of computing resources. However, the overall data of MOEA/D is the worst, with a negative optimization of 12.99%. Furthermore, although its median or average increase in resource utilization is less than that of NSGA-III-KDR, the more substantial improvement in load balancing suggests a less favorable overall evolution. Since the selection of the optimal solution using TOPSIS is random, not every iteration of MOEA/D yields a solution that outperforms those of other algorithms.
@@ VII-C Simulation of Proposed Pseudonymous Scheme
TABLE V: Comparison of computational overhead
Scheme SIPG CIMS IDMV Total
Bagga et al.’s scheme [32] $4T_{pm}^{ecc}+2T_{mtp}+1T_{ha}=12.867$ msms $3T_{pm}^{ecc}+3T_{pa}^{ecc}+3T_{ha}=1.86$ msms $3T_{bp}+4T_{pm}^{ecc}+4T_{pa}^{ecc}+3T_{ha}=6.847$ msms $21.574$ msms
Shen et al.’s scheme [33] $3T_{bp}+4T_{pm}^{bp}+1T_{pa}^{bp}+2T_{ep}+4T_{ha}=9.505$ msms $1T_{bp}+1T_{pm}^{bp}+1T_{ep}+1T_{ha}=2.745$ msms $2T_{bp}+2T_{pm}^{bp}+1T_{pa}^{bp}+1T_{ep}+1T_{ha}=5.481$ msms $17.731$ msms
Yang et al.’s scheme [34] $17T_{pm}^{ecc}+5T_{pa}^{ecc}+11T_{ha}=10.522$ msms $1T_{ha}=0.001$ msms $3T_{pm}^{ecc}+4T_{pa}^{ecc}+2T_{ha}=1.86$ msms $12.383$ msms
Wang et al.’s scheme [35] $4T_{pm}^{ecc}+1T_{pa}^{ecc}+2T_{ha}=2.475$ msms $9T_{pm}^{ecc}+4T_{pa}^{ecc}+5T_{ha}=5.571$ msms $3T_{pm}^{ecc}+2T_{pa}^{ecc}+4T_{ha}=1.86$ msms $9.906$ msms
Our proposed scheme $7T_{pm}^{ecc}+1T_{pa}^{ecc}+4T_{ha}=4.331$ msms $3T_{pm}^{ecc}+1T_{pa}^{ecc}+2T_{ha}=1.857$ msms $4T_{pm}^{ecc}+2T_{pa}^{ecc}+2T_{ha}=2.476$ msms $8.664$ msms
In this part, we evaluate the computational overhead of the proposed pseudonym scheme by comparing it with several existing pseudonymous schemes. Each operation and phase is tested 20 times separately, and the average value is recorded as the experimental data to eliminate the influence of hardware fluctuations during the operation. First, the BN curve [46] with a 128-bit security level is selected to implement the bilinear group. The execution time of basic cryptographic operations is shown in Table VI. Generally, the execution times required for bitwise XOR and modular multiplication are significantly lower compared to other cryptographic operations, and are thus disregarded. Moreover, we utilize SHA-256 as the hash function. Then, we assess the time costs associated with system initialization and pseudonym generation (SIPG), certificate issuance or message signing(CIMS), and identity or message verification (IDMV) in the pseudonym authentication process, comparing the proposed scheme with those in [32, 33, 34, 35], as shown in Table V.
TABLE VI: Execution time of cryptographic operations
@@
Next, considering the diversity of computational devices, we test the portability of the proposed pseudonym scheme. We consider five different devices, including a workstation with an Intel i9-12900k@3.9GHz and 64GB RAM (Intel i9), a computer with an i7-12700@2.1GHz and 32GB RAM (Intel i7), a computer with an i5-8500@3GHz and 8GB RAM (Intel i5), a smartphone with Snapdragon 7+ Gen 2 and 16GB RAM (Snapdragon 7+ Gen 2), and a MacBook with an M1 chip and 16GB RAM (Apple M1). We test the time consumption for each of the five phases of the pseudonym authentication process on these devices: System Initialization (SI), Computing Devices Registration (CR), Pseudonym Generation (PG), Certificate Issuance (CI), and Identity Verification (IV), including the Total Time (TT) for the entire process, as shown in Fig. 9. Time consumption varies across devices due to differences in CPU performance. Specifically, the MacBook requires the longest total time to complete a pseudonym authentication, at 20.813ms20.813\ ms. The SI is the most time-consuming, as it involves generating system keys and registering user information. However, typically, SI is only performed once, whereas CR, PG, CI, and IV take less than 8ms8\ ms across all devices, with PG taking the longest at 7.628ms7.628\ ms on the smartphone. This demonstrates the good portability and lightweight of the proposed pseudonym scheme.
@@ VIII Conclusion
This article investigates improving computing resource utilization within 6G CPN and proposes BECS, a privacy-preserving computing resource sharing mechanism. We consider various communication, computing, and service factors in 6G CPN, modeling them as a six-objective MOOP. Meanwhile, we utilize the proposed NSGA-III-KDR to find optimal solutions for this MOOP. Additionally, we introduce a novel pseudonym scheme to protect the privacy of users engaged in computing resource sharing. Extensive simulations demonstrate the effectiveness of BECS. Moving forward, we intend to further explore computing sharing in dynamic real-time scenarios and investigate advanced access control solutions to protect the task payload.
[Uncaptioned image] Kun Yan (Member, IEEE) received the B.S. degree in internet of things engineering from Xi’an University of Technology, Xi’an, China, in 2019, and the Ph.D. degree in cryptography from Xidian University, Xi’an, China, in 2025. He is currently a Postdoctoral Fellow at The Hong Kong Polytechnic University, Hong Kong. His research interests include 6G networks, resource management, internet of things, and blockchain.
[Uncaptioned image] Wenping Ma (Member, IEEE) received the B.S. and M.S. degrees in fundamental mathematics from Shaanxi Normal University, Xi’an, China, in 1987 and 1990, respectively, and the Ph.D. degree in communication and information system from Xidian University, Xi’an, in 1999, where he is currently a Full Professor with the School of Telecommunications Engineering. His current research interests include information theory, communication theory, blockchain, and 6G networks security.
[Uncaptioned image] Shaohui Sun (Senior Member, IEEE) received his Ph.D. from Xidian University, Xi’an, China, in 2003. From March 2003 to June 2006, he was a postdoctoral research fellow at the Datang Telecom Technology and Industry Group, Beijing, China. From June 2006 to December 2010, he worked at the Datang Mobile Communications Equipment Co., Ltd., Beijing. Since January 2011, he has been the Chief Technical Officer with Datang Wireless Mobile Innovation Center of the Datang Telecom Technology and Industry group. His current research interest includes advanced technologies related to B5G/6G.